Simplify the expression. Assume all variables are positive. Leave answer in radical form, no decimals. Let r(x)=2x and s(x)=x^3-3. Find s(r(2)).
Answer by rapaljer(4551)

Let r(x)=2x and s(x)=x^3-3.
Find s(r(2)).

r(2) = 2*2=4

s(r(2))
=s(4)
=4^3-3
=64-3
=61

R^2